### v2

# Unsupervised learning for inverse problems in computed tomography

Abstract
This study presents an unsupervised deep learning approach for computed tomography (CT) image reconstruction, leveraging the inherent similarities between deep neural network training and conventional iterative reconstruction methods. By incorporating forward and backward projection layers within the deep learning framework, we demonstrate the feasibility of reconstructing images from projection data without relying on ground-truth images. Our method is evaluated on the two-dimensional 2DeteCT dataset, showcasing superior performance in terms of mean squared error (MSE) and structural similarity index (SSIM) compared to traditional filtered backprojection (FBP) and maximum likelihood (ML) reconstruction techniques. Additionally, our approach significantly reduces reconstruction time, making it a promising alternative for real-time medical imaging applications. Future work will focus on extending this methodology to three-dimensional reconstructions and enhancing the adaptability of the projection geometry.
1 Introduction
Deep learning has emerged as a transformative technology in various fields, including medical imaging. In the context of CT reconstruction, different methods have been developed to enhance image quality through supervised and unsupervised training techniques. Notably, certain approaches that leverage supervised training, i.e. utilizing ground-truth images for calculating training loss, have demonstrated significant improvements in image quality. For instance, Jin et al., (2017) employed a U-Net architecture in the image domain to enhance reconstruction from sub-sampled projection data. He and Ma, (2018) utilized fully connected layers operating on sinogram data to learn the projection filter step, comparing their reconstructions to ground-truth images for effective supervised training. Similarly, Li et al., (2019) proposed a FBP-inspired network architecture designed for direct learning from sinogram to image and Ge et al., (2020) developed a network that integrates both a projection domain unit and an image domain unit, connected by a domain transform layer. Other methods rely on unsupervised deep image priors to replace or enhance steps in conventional iterative reconstruction (Zhang et al., (2020), Li et al., (2020)). To our knowledge, there exists no approach that leverages the similarities between conventional iterative reconstruction and deep neural network training, which both employ gradient-descent optimization.
1.1 Computed tomography as inverse problem
In CT, the reconstruction of imaging data can be framed as an inverse problem, where the goal is to recover the original image from its projections. Let $f:\mathbb{R}^{2}\to\mathbb{R}$ be the continuous function that represents the imaging object’s attenuation values. Assuming a circular measurement trajectory, the geometric forward projection (FP) in computed tomography can be expressed by the Radon transform
$\centering\mathrm{FP}(f)=\int_{-\infty}^{\infty}\int_{0}^{2\pi}f(x,y)\cdot\delta(x\cos\gamma+y\sin\gamma-s)\,d\gamma\,ds,\@add@centering$ (1)
where $\delta$ denotes the Dirac function, $\gamma$ the projection angle and $s$ the coordinate of the projection path through the object. This continuous form can be transferred to a linear discrete forward problem. Let $\vec{f}\in\mathbb{R}^{N}$ be the dicretized, vectorized form of the image $f$ and let $\vec{p}\in\mathbb{R}^{M}$ be its FP. Let the discrete version of the Radon transform from Eq. 1 be represented by a system matrix $A\in\mathbb{R}^{M\times N}$, that attributes a contribution of each pixel $f_{j},\,j=1,...,N,$ to each projection value $p_{i},\,i=1,...M$. Then, the system matrix defines the discrete FP $A\vec{f}=\vec{p}$ as well as the inverse problem
$\text{Find}\hskip 4.0pt\vec{f},\hskip 8.0pt\text{s.t.}\hskip 8.0ptA\vec{f}=\vec{p}.$ (2)
1.2 Iterative gradient-based CT reconstruction
Iterative reconstruction (IR) is a method that aims to progressively improve the reconstructed image $\vec{f}$ while adhering to Eq. 2. Figure 1 visualizes the general procedure. Let $f^{k}\in\mathbb{R}^{N}$ be the reconstructed tomogram at the iteration step $k\in\{0,\ldots,K\}$. In particular, let $f^{0}$ be the initial image, usually selected as a constant image with a very small positive value. To achieve a stepwise approximation of the current image $f^{k}$ to the underlying true image $f^{\star}$, a comparison is made between the measurement $p$ itself and the FP of the current image $f^{k}$, i.e. $p^{k}=Af^{k}$. Assuming a Gaussian distribution of the underlying projection data, the metric of discrepancy or loss $\mathcal{L}$ is chosen as the $\mathcal{L}^{2}$-norm
$\mathcal{L}(f^{k}|p)=\|p-p^{k}\|^{2}=\|p-Af^{k}\|^{2}.$ (3)
Due to overall differentiability of Eq. 3, the discrepancy $\mathcal{L}$ can be minimized by iterative gradient-based optimization until a predefined stopping criterion is reached, i.e.
$f^{k+1}=f^{k}+\alpha\cdot\mathcal{D}(\nabla f^{k})$ (4)
where the step width $\alpha$ and the search direction $\mathcal{D}$ is chosen by the respective optimization method, usually in dependence on the gradient $\nabla f^{k}$. Regarding implementation, it is important to ensure that the matrix $A^{T}$ satisfies the adjointness condition
$\langle A\vec{f},\vec{p}\rangle_{\mathbb{R}^{M}}=\langle\vec{f},A^{T}\vec{p}\rangle_{\mathbb{R}^{N}}\quad\forall\vec{f}\in\mathbb{R}^{N},\forall\vec{p}\in\mathbb{R}^{M},$ (5)
where $\langle\cdot,\cdot\rangle$ denotes the scalar product of the Hilbert spaces $\mathbb{R}^{N}$ and $\mathbb{R}^{M}$. This condition needs to be fulfilled for gradient-based optimization since the adjoint matrix $A^{T}$ is used as the domain transform from projection to image space - known as back projection (BP). When computing the image gradient from the projection differences, adjointness of $A^{T}$ guarantees appropriate scaling of the gradient values (Bertero et al., (2021)).
[…]
Figure 1: Gradient-based image reconstruction scheme. The colored boxes indicate the domain of the respective data ( projection domain, image domain). The boxes indicate the domain transform by FP (AA) or BP ($A^{T}$). The steps to iteratively adapt the current image $f^{k}$ with the current gradient ($\nabla$) are indicated by arrows. The loss $\mathcal{L}$ is defined by Eq. 3.
1.3 Scope of this work
Our main contribution is an unsupervised learning approach to reconstruction, treating it as an inverse problem. We leverage the similarities between this problem and gradient-based iterative reconstruction. Our focus is on exploring the possibility of training a neural network how to solve an inverse problem using only the forward model of the imaging system, without relying on a ground-truth image. The objectives for network training are:
  1. A
    To demonstrate that it is possible to invert the Radon transform (Eq. 1) using the proposed unsupervised approach on high-precision data. For this objective, forward projections of provided FBP CT slices are used.
  2. B
    To reconstruct real beam-hardening affected projection data using the proposed unsupervised approach. The raw data of an open-source dataset is utilized and image quality is compared to the respective high-quality data reconstructions.
2 Methods
2.1 Forward and backward projection layers
A critical point of novel deep learning approaches in CT is the incorporation of FP and BP operators into the deep learning framework. These operators enable forward and backward pass computations from the image domain to the projection domain and vice versa.
The deep learning framework of this work is implemented in PyTorch (Paszke et al., (2019)). We employ the projector operator definitions of the open-source library Parallelproj (Schramm and Thielemans, (2023)), which was originally designed for positron emission tomography projections utilizing PyTorch compatible tensors. The projection geometry can be defined manually to emulate a circular CT projection trajectory by explicitly defining the respective start and end points of the projection rays. Hence, a projection operator and its adjoint can be incorporated in custom FP and BP network layers. These layers now take advantage of PyTorch’s autograd function, such that no manual gradient computations are necessary. The FP and BP of Parallelproj are defined mathematically adjoint (Eq. 5), such that iterative computations of gradients are numerically consistent.
2.2 Unsupervised training scheme
[…]
Figure 2: Proposed unsupervised training scheme. The colored boxes indicate the domain of the respective data ( projection domain, image domain). The boxes indicate the domain transform by FP (AA) or BP ($A^{T}$). The steps to iteratively adapt the UNet++ weights are indicated by arrows.
Figure  2 illustrates our newly proposed unsupervised training approach. As network input, we only assume knowledge of the projection data $p$, which is transformed to the image domain by a simple BP (Radon inversion) layer $A^{T}$. The resulting blurred image $f$ is fed into a deep convolutional network, namely the U-Net++ (Zhou et al., (2018)) with the ResNeXt-101 (32 × 16d) (Xie et al., (2017)) encoder. Regarding the implementation, the Segmentation Models PyTorch (Iakubovskii, (2019)) library was employed, offering a pre-built network architecture and encoder as well as loss functions for the PyTorch machine learning framework. The network’s output is the desired high-quality image reconstruction $\hat{f}$. To enable unsupervised training, the forward pass through the network is completed with the FP layer $A$ that is applied on the image $\hat{f}$. Thus, the reconstructed data $\hat{f}$ is represented in the projection data $\hat{p}$. Hence, $\mathcal{L}_{2}$-loss calculations can be performed in the projection domain without the need for an image-domain ground truth. The backward pass propagates the calculated scalar loss value through the predicted projection tensor $\hat{p}$, the BP layer $A^{T}$ and the predicted image tensor $\hat{f}$ into the network layers. For training, the ADAM optimizer was used with a learning rate of $10^{-4}$ on an NVIDIA RTX A6000 GPU with 48 GB of memory, and no augmentations were applied to maintain physical and geometrical consistencies.
It should be highlighted here how similar the presented training approach is to gradient-based IR. Comparing Figure 1 and Figure 2, the current training prediction $\hat{f}$ coincides with the iterative scheme’s current image $f^{k}$. The IR loss and the training loss are computed with the same formula from Eq. 3.
The main difference is also the main advantage of the network approach: Instead of altering the image itself, in network training, the gradient only affects the network’s weights. This means, once trained, the network knows how to calculate the high-quality image from the blurry initial image $f$ in one step. Compared to IR, this eliminates the need for computationally extensive gradient calculations after training.
2.3 Data and reconstruction
For validation of the proposed unsupervised training approach, the publicly available 2DeteCT (Kiss et al., (2023)) dataset is employed. It is tailored for advancing machine learning techniques in CT image reconstruction, artifact reduction, and segmentation. The dataset includes over 5,000 CT reconstructed 2D FBP slices and their respective 2D projection data. The projections of each slice were acquired with three varying radiation parameter modes. We considered two of the three given modes in this work: High precision mode (90 kVp, Thoraeus filter of Sn=0.1 mm, Cu=0.2 mm, Al=0.5 mm) and beam-hardening affected mode (60 kVp, no filtration).
We employ 2800 of the 5000 provided slices and respective projection data, since a slight difference in geometry has been reported for the other measurements. This poses a critical difference for the reconstruction network since the FP and BP layers, as of now, are designed as fixed. To keep validation times during training short, we choose 100 projection datasets for validation between epochs for training supervision and early stopping. 2400 projection datasets are used for training. Approximately 10% of the available data (300 test datasets) are used for evaluation.
To compare our method, we employ FBP with a Hann filter (with cut-off at 80% of normalized maximal frequency) and gradient-based maximum likelihood (ML) reconstruction. The FBP is calculated with our in-house software based on the Reconstruction Toolkit (Rit et al., (2014)) whose projectors are parallelized on GPU. The ML optimization is performed with SciPy’s (Virtanen et al., (2020)) L-BFGS-B implementation without bound conditions or regularization. The loss and gradient computations from Eq. 3 and Eq. 4 are calculated with the adapted Parallelproj projectors (see Subsection 2.1) as well as the explicit gradient formula. We choose 100 iterations for ML optimization, which poses a good trade-off between image quality and reconstruction time in initial results: The norm of the gradients at the end of optimization calculates to an average of approx. $1.2\cdot 10^{-7}\,(\pm 4.2\cdot 10^{-8})$ and an average pixel value change of approx. 0.6% for subsequent reconstructed images, which implies sufficient convergence.
3 Results
3.1 Experiment A: Radon inversion
Experiment A was evaluated in multiple stages. Since this experiment was conducted with forward projected data for which the exact geometry was known for both forward and backward projection, we can assume perfect alignment of projected and reconstructed data with our ground truth.
As the first stage, the general behavior of our method concerning projection domain loss minimization is investigated. Figure 3 shows a qualitative example of the projection data network prediction versus the ground-truth data. The exemplary difference image as well as the quantitative MSE evaluation in Figure 4 show that our method indeed predicts projection data that is highly similar to the ground-truth projection data. Slight overestimation in the center and underestimation in the rear regions of the sinogram in Figure 3(c) could be observed. We compared our method’s results to performing FBP and gradient-descent ML on the forward projected data (details in Subsection 2.3). Our method performs significantly better than ML and FBP as calculated with a one-sided Mann-Whitney U test, which results in $p\lll 0.05$ in both cases.
Refer to caption(a) Ours(b) GT(c) Ours - GT
Figure 3: Exemplary visualization of (a) predicted projection data, (b) ground-truth projection data, and (c) difference.
Figure 4: MSE distribution of 300 projection predictions from test dataset w.r.t. ground-truth projection data. [MSE] = $1/\mathrm{cm}^{2}$
Figure 5: Quantitative results in image domain for radon inversion task on 300 test images. All methods are calculated based on forward projected data of ground-truth image. Deviations are calculated with respect to said ground-truth image. [MSE] = $1/\mathrm{cm}^{2}$
Second, the quantitative deviation of the reconstructed images to the ground-truth images was evaluated. The results for MSE and SSIM are presented in Figure 5. The distributions of MSE and SSIM values across methods appear similarly shaped. It could be observed that the images reconstructed by our unsupervised network show lower MSE and higher SSIM values than those obtained with alternative methods. Since, to our knowledge, no other unsupervised deep-learning methods exist, no comparisons to other methods from this field were performed.
[…]
Figure 6: Reconstructed images for different methods. The image was chosen as one of the worst MSE values (MSE = $6.8\cdot 10^{-7}\,\frac{1}{\mathrm{cm}^{2}}$) between our method’s result and the ground-truth (GT) image. First row in (a) is shown in level/window=0.03/0.06 $\frac{1}{\mathrm{cm}}$. The orange rectangle indicates the ROI for Figure (b). First row in (b) shown in level/window=0.05/0.1 $\frac{1}{\mathrm{cm}}$.
Third, the reconstructed images were evaluated qualitatively. In Figure 6, a dataset of high deviation to the ground-truth image was evaluated. The figure shows the comparison across methods and the corresponding difference images for a whole reconstructed slice as well as an enlarged region for further inspection. High image quality can be observed in all reconstructions. The difference images reveal that while FBP and ML exhibit a typical noise level all over the reconstructed image, our predicted image mostly shows differences in the higher density regions.
To further investigate the noise properties in these images, Figure 7 shows the distribution for an exemplary region that was assumed to be of a constant image value. Here, more deviation from the ground-truth mean value (marked as a vertical line) indicates a higher standard deviation and hence a higher noise level. As highlighted by the shadowed regions, our approach reduced the standard deviation while adhering closely to the correct mean value. While not reproducing the ground-truth image exactly, this behavior suggests that the network produces smoother imaging results than FBP and ML.
Refer to captionRefer to caption
Figure 7: Image value distribution of an exemplary constant image region for the different methods. The region of interest (ROI) is marked in red. Gaussian distribution curves are fitted to emphasize distribution differences. The black vertical line indicates the mean value of the ground-truth image in the ROI. The orange vertical line indicates our prediction’s mean value. The shadowed regions indicate the standard deviation of the GT ( ) and our method ( ).
3.2 Experiment B: Reconstruction of beam-hardening afflicted data
In Experiment B, real beam-hardening affected projection data was used for reconstruction. Again, the projection domain difference to the ground-truth data was evaluated. Figure 8 shows the distribution of MSE of the test datasets for all three considered methods. It can be observed that, in terms of similarity to raw data, FBP and ML perform slightly worse than the network reconstruction. It should be noted here that the reconstruction for our method is pure network inference. For a number of 10 reconstructions, we observe a mean reconstruction time of 438 ms ($\pm$ 51 ms) per dataset for the network reconstruction, whereas FBP takes 1717 ms ($\pm$ 47 ms) and ML reconstruction takes approx. 63 400 ms ($\pm$ 2 200 ms) on average.
Figure 8: Evaluation of Experiment B in projection domain. MSE distribution (w.r.t. ground-truth raw data) of 300 forward projected images from test dataset for different reconstruction methods.
The given image slices from the 2DeteCT dataset using FBP merely pose one possible reconstruction result. Thus, there exists no true ground-truth image information in Experiment B and the quantitative MSE evaluation in the image domain should be omitted. Instead, we can qualitatively compare performance by evaluating FBP reconstructions from high-quality raw data mode (as was used in Experiment A). We indicate this ”ground-truth” image with the abbreviation GT. Figure 9 shows the qualitative comparison in the image domain. Since the images look very similar, difference images to our method highlight their differences. The deviations seem to lie mostly in edges, higher density as well as artifact regions. The orange arrows in Figure 9(b) indicate that some details might get lost in our approach due to smoothing - which was observable in the noise analysis of Experiment A as well. In summary, our method performs subjectively similarly in the image domain while quantitatively outperforming in the projection domain and needing a fraction of the reconstruction time of conventional ML.
[…]
Figure 9: Reconstructed images for different methods. The ground-truth (GT) image is generated from high-quality acquisition mode projection data without beam-hardening. Difference images are shown w.r.t. our method. The first row images are displayed in level/window=0.05/0.1 $\frac{1}{\mathrm{cm}}$. The orange rectangle indicates the ROI for Figure (b). First row shown of Figure (b) in level/window=0.05/0.1 $\frac{1}{\mathrm{cm}}$.
4 Discussion
Does our approach perform better than FBP and ML in Experiment A? Yes. The ground-truth image in this case is the mathematically optimal solution for the pure Radon inversion task. As can be observed in Figure 5 and Figure 6 the network predicts reconstructed images that are closer to the ground-truth than FBP and ML. However, it is noticeable that uncertainty arises mainly in higher density material regions.
For high-quality data, as provided in Experiment A, FBP outperforms ML. We assume that this is due to the high quality of the underlying data, such that FBP performs well. Additionally, the stopping criterion for ML convergence has been configured such that optimization time does not become unreasonably long. It is expected that a higher number of iterations will outperform FBP in this case. As can be observed in the real-data case (Experiment B), ML does provide better results with the same tolerance choices and number of iterations for real beam-hardening affected data.
The noise analysis in Figure 7 shows that our method tends to smooth images more than FBP and ML. In total, in this constructed scenario with perfect knowledge of geometry and expected image values, the unsupervised network outperforms traditional methods.
How does the approach perform in Experiment B? Figure 8 shows that adherence to the raw projection data is better for our reconstruction network than FBP and ML. A detailed investigation of the MSE distribution showed that especially background noise was reduced by our approach. When inspecting images in Figure 9, some details seem lost in reconstruction. It is possible that the choice of the $\mathcal{L}_{2}$ loss leads to this behavior, since it can lead to over-smoothing of the image in combination with network training. It is to be investigated where this difference to ML reconstruction stems from, considering that ML also utilizes $\mathcal{L}_{2}$ loss. Insufficient fine-tuning of the network could also be a possible reason for this behavior.
Overall, our approach performs similarly to the established methods while reducing reconstruction time significantly. It should be mentioned that the network can still be improved by additional samples in an online training approach. Of course, the reconstructed images from our approach could also be used as initial images for conventional iterative reconstruction to reduce the number of iteration steps.
4.1 Limitations
The proposed approach presents several notable limitations. Firstly, this work is confined to two-dimensional reconstructions. For three-dimensional scenarios, substantial GPU memory will be necessary, which results in increased hardware requirements for effective implementation. The approach also operates within a fixed geometric framework, which limits the amount of applicable data immensely. Additionally, in this work, no data augmentations were applied. These usually increase the network’s ability to generalize across different datasets. While the training process is time-consuming, the application of the trained network for reconstruction is executed with notable speed. These limitations indicate areas for potential improvement and further research to enhance the overall efficacy of the method.
4.2 Future work
Based on the observed limitations, we propose the following advancements for our future work. To expand to 3D reconstruction, the approach should be tested with cone-beam CT datasets for which raw data is available. A detailed description of the projection geometry is needed for such a dataset as well. Additionally, high variability of the data and the availability of augmentations will help achieve a stable and high-performing network.
The geometry of the projection layers should be adaptable to the dataset, which will increase the amount of suitable data. This can easily be realized but needs either additional computations when switching datasets or larger memory for different pre-calculated projection layers. These additional resources could be realized with multi-GPU parallelization, which was out of the scope of this work.
Overall, the inclusion of unsupervised neural networks into image reconstruction creates new opportunities for image manipulation and use-case adaptation. This holds especially in comparison to conventional iterative methods, where regularizing terms and image gradients have to be designed carefully and computed expensively. While, in this work, we focused on the specific problem of CT imaging, the general approach of unsupervised training can be transferred to arbitrary inverse problems for which a continuous formulation (i.e. Eq. 2) of the forward problem can be modeled.